is 2y=5x+6
-10+4y=8 independent, depehndent, inconsistent or consisten

Answer by stanbon(75887)

is
2y=5x+6
-10x+4y=8
--------------
y = (5/2)x + 3
y = (5/2)x + 2
----
You have 2 parallel lines.
The system is inconsistent.
----------------
Cheers,
Stan H.
================
independent, depependent, inconsistent or consistent.
================
Cheers,
Stan H.
================